You are the CEO of a private company and want to buy it. There is one...

  1. You are the CEO of a private company and want to buy it. There is one major shareholder who owns 70 percent of the firm. The remaining 30 percent is held by oceanic shareholders. Based on a DCF model, you estimate the base value of the firm to be $100 million. This excludes any liquidity discount and/or control premium. Based on your observations of the market, liquidity discounts are about 30 percent and control premiums are about 15 percent for similar deals. Your firm has 80 million shares outstanding.

What would you offer to pay in total and per share to the controlling shareholder? To the oceanic shareholders?

What if it were not a private company but a public company. Show how would that change your calculation and answers?
Answer #1

I have answered the question below.

Please upvote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

When it is a Private company A Base value of firm ($ mil] B Liquidity discount adjustment factor (1-0.30] C Control Premium a

Working for Per share calculation:

In the first case, it is found out that the value of firm is 80.50Mn

Also given that there are 80 Mn shares outstanding,

=> Value/share = 80.5/80 = 1.01 (apprx)

Even though the value changes between 56.35 Mn and 24.15 Mn, the per share value remains unless mentioned that it is acquired at premium ( not mentioned anywhere). So, the value/share remains the same as $1.01 for both companies

In the second case, it is found out that the value of firm is 115 Mn

Also given that there are 80 Mn shares outstanding,

=> Value/share = 115/80 = 1.44 (apprx)

Even though the value changes between 80.5 Mn and 34.5 Mn, the per share value remains unless mentioned that it is acquired at premium ( not mentioned anywhere). So, the value/share remains the same as $1.44 for both companies